IF THE MAXIMUM & MINIMUM VALUE OF -


y= x2-3x+c / x2+3x+c  ARE 7 AND 1/7 ,THEN VALUE OF c IS ...?

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear rahul

y= x2-3x+c / x2+3x+c

cross multiply and simplify

x2(y-1) +x(3y+3) +c(y-1)=0

this equation has a real value so discrimnent will be ≥ o

 (3y+3)2 -4c(y-1)(y-1)≥0

simplyfy

{y(3+2√c) -(2√c-3)}{y(2√c -3) -(3+2√c)}≤0

            (2√c-3)/  (3+2√c) ≤  y≤(3+2√c)/(2√c -3)

compair with given condition

(3+2√c)/(2√c -3) =7

 c= 9/4

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin